# Improving Autoregressive Visual Generation with Cluster-Oriented Token Prediction

### v1

@@ 2.2 Visual Generation Models
Continuous-valued Visual Generation. Generative Adversarial Networks (GANs) [16, 25, 14, 43], consisting of a generator and discriminator, employ adversarial training to learn the target distribution. Diffusion Models [20, 35, 15, 13, 21, 31] iteratively add and remove noise to approximate the real data distribution, using a Markov process for the forward path and learning the reverse denoising process. These models have achieved significant success in image generation and have been extended to video generation [17, 4, 42], enhancing video quality. However, both GANs and Diffusion Models, which model continuous distributions, face challenges when integrating with discretely modeled LLMs. Thus, exploring discrete modeling methods for image generation is necessary.
@@ 4.3 Ablation Study
Ablation on codebook rearrangement and $\mathcal{L}_{CCE}$. We first evaluate the effectiveness of the codebook rearrangement strategy and the cluster-oriented cross-entropy loss $\mathcal{L}_{CCE}$ We conduct experiments on three ablated models: the model without both codebook rearrangement and $\mathcal{L}_{CCE}$ (original LlamaGen), the model without codebook rearrangement, and the model without $\mathcal{L}_{CCE}$. The results are shown in Tab 3. It can be seen that the model with only the codebook rearrangement (w.o. CCE loss) performs almost the same as LLamaGen. This is because the original token-oriented cross-entropy loss does not consider cluster information and makes no distinction between embeddings with distant indices and those with close indices, thus failing to utilize the similarity within local clusters. Therefore, even with codebook rearrangement, the result remains almost the same as the original LlamaGen. For the model with CCE loss only (w.o. rearrangement), we treat adjacent embeddings in a window (e.g., $0\text{-}128$-th codes, $128\text{-}256$-th codes) as a “cluster”. Without the clustering and reorderng process, codes in such “clusters” are not similar to each other, therefore CCE loss cannot function well and only slightly improves the performance. In contrast, our model with both codebook rearrangement and $\mathcal{L}_{CCE}$ has the best performance in fitting the ImageNet dataset.
@@ 5 Conclusion
In this paper, we analyzed the difference between the natural language and images in LLM-based visual generation. We find that similar image embeddings in the codebook can produce similar images. Based on this observation, we propose IAR, an Improved Autoregressive Visual Generation Method that effectively enhances the training efficiency and quality of the LLM-based visual generation model. We first conduct Codebook Rearrangement with a balanced K-means clustering algorithm to reorder the codebook into clusters with equal sizes, where the image embeddings within each cluster share similarities. Then, we introduce a Cluster-oriented Cross-entropy Loss to enable the model to learn the target cluster, which can guarantee a good generation quality even if the model predicts a wrong image token. We validate the effectiveness of our IAR on LlamaGen, and find that IAR can largely improve the training efficiency and generation quality across different parameter scales. Moreover, our IAR can be seamlessly applied to various LLM-based visual generation model, pointing out a new direction to improve the LLM-based visual generation field.
@@ Appendix A Overview
  • More visualization results (Sec. F);
  • Future work (Sec. G).
We also provide the code of our model in the supplementary material, which can be easily applied to the existing autoregressive models.
@@ B L XL XXL B L XL XXL
Appendix F More Visualization Results
@@
Appendix G Future Work

### v2

@@ 2.2 Visual Generation Models
Continuous-valued Visual Generation. Generative Adversarial Networks (GANs) [16, 27, 14, 47, 46], consisting of a generator and discriminator, employ adversarial training to learn the target distribution. Diffusion Models [20, 38, 15, 13, 23, 21, 34] iteratively add and remove noise to approximate the real data distribution, using a Markov process for the forward path and learning the reverse denoising process. These models have achieved significant success in image generation and have been extended to video generation [17, 4, 45, 24], enhancing video quality. However, both GANs and Diffusion Models, which model continuous distributions, face challenges when integrating with discretely modeled LLMs. Thus, exploring discrete modeling methods for image generation is necessary.
@@ 4.3 Ablation Study
Ablation on codebook rearrangement and $\mathcal{L}_{CCE}$. We first evaluate the effectiveness of the codebook rearrangement strategy and the cluster-oriented cross-entropy loss $\mathcal{L}_{CCE}$ We conduct experiments on three ablated models: the model without both codebook rearrangement and $\mathcal{L}_{CCE}$ (original LlamaGen), the model without codebook rearrangement, and the model without $\mathcal{L}_{CCE}$. In contrast, our model with both codebook rearrangement and $\mathcal{L}_{CCE}$ has the best performance in fitting the ImageNet dataset.
@@ 5 Conclusion
In this paper, we analyzed the difference between the natural language and images in LLM-based visual generation. We find that similar image embeddings in the codebook can produce similar images. Based on this observation, we propose IAR, an Improved Autoregressive Visual Generation Method that effectively enhances the training efficiency and quality of the LLM-based visual generation model. We first conduct Codebook Rearrangement with a balanced K-means clustering algorithm to reorder the codebook into clusters with equal sizes, where the image embeddings within each cluster share similarities. Then, we introduce a Cluster-oriented Cross-entropy Loss to enable the model to learn the target cluster, which can guarantee a good generation quality even if the model predicts a wrong image token. We validate the effectiveness of our IAR on LlamaGen, and find that IAR can largely improve the training efficiency and generation quality across different parameter scales. Moreover, our IAR can be applied to various LLM-based visual generation models, pointing out a new direction to improve the LLM-based visual generation field.
Acknowledgments
This work was supported by Shanghai Sailing Program (22YF1420300), National Natural Science Foundation of China (No. 62302297, 72192821, 62272447, 62472282, 62472285), Young Elite Scientists Sponsorship Program by CAST (2022QNRC001), the Fundamental Research Funds for the Central Universities (YG2023QNB17, YG2024QNA44), Beijing Natural Science Foundation (L222117).
@@ Appendix A Overview
  • Analysis on the cluster-oriented cross-entropy loss $\mathcal{L}_{CCE}$ (Sec. F);
  • Analysis on the influence of code-rearrangement quality (Sec. G);
  • Experiments on different VQVAEs (Sec. H);
  • More visualization results (Sec. I);
  • Future work (Sec. J).
@@ B L XL XXL B L XL XXL
Appendix F Analysis on the Cluster-oriented Cross-entropy Loss
In Tab. 3 of the main paper, the model trained with only the cluster-oriented cross-entropy loss $\mathcal{L}_{CCE}$ (without codebook rearrangement) also improves generation performance (FID 6.96 vs. 7.14 for the baseline). This improvement arises because $\mathcal{L}_{CCE}$ enhances the probability of predicting the correct cluster, which is computed based on the probabilities of all tokens within the cluster. Minimizing $\mathcal{L}_{CCE}$ consequently increases the probability of the target token. Furthermore, since $\mathcal{L}_{CCE}$ boosts the probabilities of all tokens in the target cluster, it can be viewed as a variant of label smoothing, which is known to improve generalization and model calibration in classification networks [33]. However, as shown in Table A7, standard label smoothing is not well-suited for autoregressive visual generation models, where token prediction accuracy is typically low (e.g., 2%–4% top-1 accuracy, as shown in Table A6), in contrast to traditional classification tasks with significantly higher accuracy (e.g., $>70\%$). In contrast, $\mathcal{L}_{CCE}$ performs structured smoothing within specific ranges rather than uniformly smoothing all tokens, leading to improved AR generation quality. While $\mathcal{L}_{CCE}$ enhances model performance, the best results are achieved only when it is combined with our codebook rearrangement strategy. Therefore, the good performance of our IAR is attributed to both the codebook rearrangement strategy and the cluster-oriented cross-entropy loss.
@@ Appendix F Analysis on the Cluster-oriented Cross-entropy Loss
LlamaGen 6.05 182.5 0.84 0.42
LlamaGen + IAR 5.77 192.5 0.85 0.42
LlamaGen + LS 8.58 166.81 0.80 0.40
Table A7: Comparison with the model with label smoothing, where we train LlamaGen-B with default smoothing factor 0.1 for 300 epochs. It shows that the vanilla label smoothing is not suitable for AR generation models with low token prediction accuracy.
Appendix G Analysis on the Influence of Code-rearrangement Quality
Our IAR employs a code-rearrangement strategy to cluster similar codes, with the hypothesis that better clustering quality theoretically enhances generation performance. To validate this, we average the mean distance of each cluster as a metric to evaluate clustering quality, where a lower mean distance indicates a higher codebook rearrangement quality. We train LlamaGen-B on codebooks with varying clustering qualities (different mean distances, obtained by setting different clustering iterations in Balanced k-means Clustering) for 100 epochs to investigate how clustering quality affects performance. Results in Table A8 demonstrate that improved clustering quality leads to better model generation performance.
[…]
0.689 6.77 171.73 0.84 0.42
0.850 6.85 170.54 0.84 0.41
1.239 6.91 169.32 0.84 0.41
Table A8: Generation performance from codebooks with different clustering quality.
Appendix H Experiments on Different VQVAEs
To show the effectiveness and generalization ability of our method, we further conduct experiments on the VQVAE from VQGAN [14], where we train LlamaGen-B on it for 100 epochs. As shown in Tab. A9, our IAR consistently improves the generation quality for the model based on VQGAN, demonstrating the effectiveness of our IAR across different VQVAEs.
[…]
LlamaGen 7.14 166.38 0.84 0.40
LlamaGen 6.77 171.73 0.84 0.42
VQGAN 6.90 176.71 0.84 0.40
VQGAN 6.75 194.64 0.85 0.40
Table A9: Experiments on different VQVAEs. Our IAR consistently improves the performance of the model trained on different VQVQEs.
[…]
LlamaGen 2.06 0.20 4.30 20.79
LlamaGen 0.69 0.18 1.91 20.79
VQGAN 24.75 5e-7 1072.06 20.00
VQGAN 8.69 5e-7 44.68 20.00
Table A10: Comparison between the original and rearranged codebooks from Llamagen and VQGAN. Our code rearrangement strategy can consistently improve the inner-cluster similarity for both the two codebooks.
We further compute the mean / closest / largest $\mathcal{L}_{2}$ distance in each cluster for the rearranged and original codebook from LlamaGen and VQGAN, to validate the effectiveness of our code rearrangement strategy. The results in Tab. A10 show that our method stably decreases the three distances. Moreover, it can be seen that the closest distance in VQGAN is $5\times 10^{-7}$, indicating that some embeddings are almost the same, resulting in some embeddings being wasted (e.g., the closest distance of VQGAN (5e-7) is very low, causing the worst reconstruction PSNR (20.00)). Therefore, it motivates us that the closest distance can be used as a metric to evaluate the training of a VQVAE.
Appendix I More Visualization Results
[…]
Appendix J Future Work